OpenStudy (anonymous):

FInd the similarity ratio of a cube with volume 216 ft^3 to a cube with colume 1000 ft ^3 Please Help

OpenStudy (jdoe0001):

if something is say similar to something else by some ratio the ratio will be a/b corresponding units a box is 3, a similar box is 1, what's the ratio the bigger box is 3/1 times the small one the smaller box is 1/3 times the big one

OpenStudy (anonymous):

the answers are A3:5 B36:100 C 100:36 D 5:3 @jdoe0001

OpenStudy (anonymous):

so its either A or B? @jdoe0001

OpenStudy (jdoe0001):

hmm sorry, was a bit ....

OpenStudy (jdoe0001):

$$\bf 216 = 6^3\\ 1000 = 10^3\\ \text{the ratio is 6:10 or } \cfrac{6}{10} \implies \cfrac{3}{5} $$

OpenStudy (jdoe0001):

or 3:5 or 5:3 depending on how you arrange them